OpenStudy (anonymous):

Evaluate the definite integral 7^3xdx from [0,1]

OpenStudy (abb0t):

\[\int\limits_{0}^{1} 7^{3x}dx\]???

OpenStudy (anonymous):

yes

OpenStudy (abb0t):

I believe it's: \[\int\limits_{a}^{b}a^tdt = \frac{ a^t }{ \ln(t) }+C\]

OpenStudy (abb0t):

oops, ignore the constant, obviously, this is a definite integral and there is no constant.

OpenStudy (anonymous):

right lol ok I'll try it. thanks